OpenStudy (anonymous):

Katherine has $140 in the bank and is saving $6 per week. Abbie has $462 in the bank, but is spending at a rate of $10 per week. Which equation will determine when they will have the same amount of money in the bank? A) 140 + 6x = 462 + 10x B) 140 + 6x = 462 - 10x C) 140 + 10x = 462 - 6x Eliminate D) 140 - 6x = 462 + 10x

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

need little help

OpenStudy (gorv):

let they saved for x week ....we dont now the excat number of week so considering it as x

OpenStudy (anonymous):

ok

OpenStudy (gorv):

kath had 140 already + save 6 per week

OpenStudy (gorv):

so in x week total saving = ??

OpenStudy (anonymous):

146

OpenStudy (anonymous):

136

OpenStudy (gorv):

no..in 1 week she saves 6 in x week she will save ??

OpenStudy (gorv):

=6*x

OpenStudy (anonymous):

i dont get this

OpenStudy (gorv):

like price of one chocolate is 5 so price of 10 chocolate will be = ??

OpenStudy (anonymous):

50$

OpenStudy (gorv):

what if we replace 10 by x???

OpenStudy (gorv):

for x chocolate ??

OpenStudy (anonymous):

we wont know

OpenStudy (gorv):

why ?? wont it be 5*x??

OpenStudy (gorv):

where x is no of chocolate

OpenStudy (gorv):

for one chocolate x=1 so total=5 for 2 choco x=2 so total= 5*2

OpenStudy (gorv):

for 10 choco x=10 total=5*10

OpenStudy (anonymous):

oh

OpenStudy (gorv):

what u say ??

OpenStudy (gorv):

like she saves 6 in one week ....so in x week ??

OpenStudy (anonymous):

a

OpenStudy (gorv):

no she is spending so we will subtract ...his saving is decreasing

OpenStudy (gorv):

-10x

OpenStudy (anonymous):

B im right

OpenStudy (anonymous):

its B

OpenStudy (gorv):

yeah u r absolutely right mate :)

OpenStudy (anonymous):

tyvm

OpenStudy (gorv):

wlcm:)

OpenStudy (anonymous):

can you helkp me one 1 more?

OpenStudy (anonymous):

help*

OpenStudy (gorv):

post new Q

OpenStudy (anonymous):

ok